# Decomposing Representation Space into Interpretable Subspaces with Unsupervised Learning

### v1

@@ C.2 Toy Model Subspace Partitioning
For toy setting, we use a slightly different version from Algorithm 1. Specifically, we find it is not necessary to search over a huge amount of subspace activations, so we search for the nearest neighbor inside the batch (Eq. 4) thus simplify the algorithm. In addition, we find $\mathbf{R}$ usually get stuck at sub-optimal point. When observing $\mathbf{R}\mathbf{W}$, we usually find one subspace is computed from two feature group, and this is usually accompanied with higher overall training loss, and more training steps and changing hyperparameters in Adam optimizer do not seem to help. We apply the following heuristics: we take the subspace with the largest loss and measure MI between this and all other subspaces, and multiply the MI values with loss of all other subspaces, then divide the resulting values (always $>$ 0) by their sum as the probability by which we randomly choose the other subspace. This process selects two large loss subspaces which encodes similar information. We multiply their corresponding rows in $\mathbf{R}$ with a random orthogonal matrix ($\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$), thus re-initialize the weights for these two subspaces by mixing these two subspaces, while the whole matrix $\mathbf{R}$ remains orthogonal. We find that this technique helps $R$ achieve lower training loss and produce a more ideal $\mathbf{R}\mathbf{W}$. In other words, we have a better chance to find global minimum. But we also find this becomes less effective when number of features and dimension become larger. This also calls for more study on this atypical optimization problem, i.e., with orthogonality constraint and ignorance of distance change to all but the nearest neighbor when computing gradients.
@@
We train $\mathbf{R}$ for toy models whose orthogonal structure is clear. Concretely, we present 4 experiments here, corresponding to toy models visualized in Fig. 1(a), Fig. 7(b), Fig. 8(a) and Fig. 8(b). The results are shown in Fig. 9, 10, 11, 12. We can see the proposed method is quite effective. Before merging, each subspace gradually learns to encode only one feature group. For example, in Fig. 9c, each 2D subspace in $\boldsymbol{\hat{h}}$ is computed either from the first or the second feature group (e.g., the first 2 rows only have nonzero entries in first 20 columns, corresponding to the first group). In Fig. 10, at the beginning the third subspace encodes all 4 groups (a), then after 3k steps it only encodes the 1st and 3rd group (b), then later it encodes only the 1st group. We can clearly see how NDM training objective leads $\mathbf{R}$ to evolve towards the desirable direction. Loss values also confirm the alignment between neighbor distances and “purity” of subspace (encoding one group instead of a mixture of groups). We denote the loss when search number is 4 as $l_{N=4}$, and similarly $l_{N=1024}$. We take the mean of loss across subspaces instead of sum when reporting them. For example, Fig. 9b’s loss values are: $l_{N=4}=0.212$ and $l_{N=1024}=0.011$. After more training when subspaces are “pure” (subfigure c), loss values are: $l_{N=4}=0.194$ and $l_{N=1024}=0.004$. Another example, Fig. 10 loss values are: $l_{N=4}=0.340$ and $l_{N=1024}=0.016$ for subfigure b, and $l_{N=4}=0.322$ and $l_{N=1024}=0.004$ for subfigure c. Importantly, we also see that $\mathbf{R}$ sometimes does not find global minimum. In Fig. 11 and 12, we see some sub-optimal subspaces, such as the 3rd subspace from the last in Fig. 11c, which is the reason for later imperfection of final $\mathbf{R}$. Moreover, we also see that the proposed MI-based merging works well. It correctly merges the subspaces encoding the same group. For example, in Fig. 10, normalized MI between 1st and 5th subspace is 0.55, and between 2nd and 6th is 0.49, normalized MI between all other pairs is below 0.03. In Fig. 9 and 10, we get the correct number of subspace and their dimension. In Fig. 11 and 12, the final configuration is imperfect but close to the ground truth. As mentioned before, we generally find the imperfection results from $\mathbf{R}$ being unable to find global minimum or disentangle feature groups for a few subspaces, the MI estimate usually works well.
@@
In summary, we find the training objective of NDM indeed encourages subspaces to encode only one feature group. Together with MI based merging, the final subspace partition is close to the groundtruth, with each subspace encoding one feature group. While training objective aligns well with our goal, the orthogonal matrix sometimes gets stuck at local minimum of this objective, resulting in sub-optimal partition. Nevertheless, it might not be necessary to achieve perfect partition to make itself useful in practice. In subfigure b of Fig. 9-12, we observe that these sub-optimal points already have a good deal of disentanglement, which could potentially be very useful when using these subspaces as analyzing unit instead of random subspaces.
@@ D.2 GPT-2 Small Test Suite
The Indirect Object Identification (IOI) task consists of sentences like: “When Mary and John went to the store, John gave a drink to _”, which should be completed with “Mary”. Wang et al. [2022] found a circuit in GPT-2 Small responsible for solving this task (Fig. 13). Besides computational graph, they also reveal the information moved between the nodes of the graph.
Figure 13: IOI circuit, Fig. 2 from [Wang et al., 2022]. Attention heads in this figure is denoted as layer_index.head_index. E.g., The second previous token head, 4.11, refers to layer 4 and attention head 11 in that layer. Note that indices start from 0.
To understand our intervention, we briefly describe the information flow inside the circuit: At token S1+1 (see Fig. 13 for the notation), the model attends the previous token and moves the name from it, which will be used as key in later induction heads. Right after layer 4, the model finishes moving previous token information, and this information is not yet used by downstream components. At token S2, the induction heads attend to S1+1, moving the relative position of S1 (i.e., it occurs before or after IO, or equivalently, S1+1 is “and” or “went”). Meanwhile, Duplicate token heads attend to S1 from S2, moving S name and relative position of S1. Therefore, at S2’s residual stream, right after layer 6 the model finishes moving position information of S1, and it is not yet moved by any downstream component. Then at END token, S-Inhibition heads attend to S2 and move the position information of S1 and as well as the S name. Right after layer 8, the model finishes moving position information of S1 and information of S name, and they are not yet used by any name mover heads. Lastly, name mover heads avoid attending S1 by using those information and attend to IO to move IO name. Note that the claims about the moved information is supported by strong causal evidence from [Wang et al., 2022], and is confirmed again with a different method in [Huang et al., 2024b].
@@
The Greater-than task [Hanna et al., 2023] consists of sentences in the form “The <noun> lasted from the year XXYY to the year XX”, such as “The war lasted from the year 1732 to the year 17”, the model should assign higher probability to years greater than 32, thus performing the simple greater-than math operation. Hanna et al. [2023] revealed a circuit in GPT-2 Small responsible for solving this task. Briefly speaking, some attention heads between layer 5 to layer 9 (both inclusive) move YY to the last token’s residual stream, and MLP layer 8 to 11 (both inclusive) specify which years are greater than YY. Therefore, right after layer 9, the model finishes moving YY to the last token, so we intervene at $\boldsymbol{h}_{\text{END}}^{9,\text{post}}$ to change the YY, where END denotes the last token.
@@ D.3 NDM Training Details on GPT-2 Test Suite
[…]
@@ D.4 Representing Patching Effect Distribution with Gini Coefficient
$\text{Var}_{s}$ (×$\times 10^{3}$) 2.1 2.7 1.3 0.6 0.8 1.4 1.7 1.0 1.0 0.4 4.0 4.8 0.3
$\Delta P_{s}$ (×$\times 10^{-2}$) 5.4 0.7 1.8 1.9 29.7 0.6 0.6 0.3 0.6 0.8 1.2 0.7 0.7
$\frac{\Delta P_{s}}{d_{s}}$ (×$\times 10^{-4}$) 4.2 0.8 1.9 3.0 46.4 1.0 0.9 0.9 2.0 2.4 3.9 2.0 2.3
$\frac{\Delta P_{s}}{\text{Var}_{s}}$ (×$\times 10^{-5}$) 2.6 0.3 1.4 2.9 39.6 0.5 0.3 0.3 0.6 2.2 0.3 0.1 2.2
@@
$\text{Var}_{s}$ (×$\times 10^{3}$) 0.3 0.3 0.2 0.2 0.3 0.3 0.4 0.3 0.3 0.4 10.7
$\Delta LD_{s}$ (×$\times 10^{-1}$) 14.8 11.6 7.6 4.2 3.1 2.5 2.9 1.4 1.6 0.9 2.2
$\frac{\Delta LD_{s}}{d_{s}}$ (×$\times 10^{-3}$) 7.7 9.1 11.9 6.6 4.8 3.9 4.6 4.4 5.0 2.9 6.8
$\frac{\Delta LD_{s}}{\text{Var}_{s}}$ (×$\times 10^{-4}$) 49.5 37.6 38.5 18.1 11.0 7.5 7.1 5.5 5.2 2.2 0.2
@@
$\text{Var}_{s}$ (×$\times 10^{3}$) 0.7 1.0 0.5 7.9 0.4 0.2 0.2 0.1 0.1 0.1 0.1
$\Delta LD_{s}$ (×$\times 10^{-1}$) 3.9 3.7 2.3 2.5 1.6 1.7 1.7 0.7 0.8 0.8 0.7
$\frac{\Delta LD_{s}}{d_{s}}$ (×$\times 10^{-3}$) 3.0 2.9 2.4 2.6 2.4 2.6 2.7 2.2 2.4 2.4 2.2
$\frac{\Delta LD_{s}}{\text{Var}_{s}}$ (×$\times 10^{-4}$) 5.5 3.6 4.7 0.3 4.0 8.1 9.1 5.2 8.0 7.7 7.0
@@ D.5 Ablation Study on GPT-2 Test Suite
search number (×$\times 2^{14}$) merging threshold merging interval avg coefficient
1 0.02 4000 0.55
1 0.02 8000 0.64
1 0.03 4000 0.66
1 0.03 8000 0.61
1 0.04 4000 0.57
1 0.04 8000 0.63
1 0.05 4000 0.52
1 0.05 8000 0.58
5 0.02 4000 0.58
5 0.02 8000 0.63
5 0.03 4000 0.64
5 0.03 8000 0.64
5 0.04 4000 0.65
5 0.04 8000 0.65
5 0.05 4000 0.61
5 0.05 8000 0.63
25 0.02 4000 0.57
25 0.02 8000 0.66
25 0.03 4000 0.63
25 0.03 8000 0.65
25 0.04 4000 0.68
25 0.04 8000 0.71
25 0.05 4000 0.63
25 0.05 8000 0.71
Table 5: Combinations of different search number, merging threshold, and merging interval, and the resulting average Gini coefficient on GPT-2 test suite. The one shown in Table 1 is the third row from the last, which is slightly better than the last row if we show more precision. Other hyperparameters are fixed, and described in D.3.
@@
search number (×$\times 2^{14}$) merging threshold avg coefficient
@@
search number (×$\times 2^{14}$) merge threshold avg coefficient
@@ D.6 Knowledge Conflict Experiment
[…]
@@
After training the estimated MI decreases significantly, see Fig. 16 and 17. Interestingly, as we can see, the original basis of the representation space of Qwen2.5-1.5B is quite special, simply partition on standard basis can produce many independent subspaces.
@@
We show baseline results in Fig. 18 and 19.

### v2

@@ C.2 Toy Model Subspace Partitioning
For toy setting, we use a slightly different version from Algorithm 1. Specifically, we find it is not necessary to search over a huge amount of subspace activations, so we search for the nearest neighbor inside the batch (Eq. 3) thus simplify the algorithm. In addition, we find $\mathbf{R}$ usually get stuck at sub-optimal point. When observing $\mathbf{R}\mathbf{W}$, we usually find one subspace is computed from two feature group, and this is usually accompanied with higher overall training loss, and more training steps and changing hyperparameters in Adam optimizer do not seem to help. We apply the following heuristics: we take the subspace with the largest loss and measure MI between this and all other subspaces, and multiply the MI values with loss of all other subspaces, then divide the resulting values (always >> 0) by their sum as the probability by which we randomly choose the other subspace. This process selects two large loss subspaces which encodes similar information. We multiply their corresponding rows in $\mathbf{R}$ with a random orthogonal matrix ($\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$), thus re-initialize the weights for these two subspaces by mixing these two subspaces, while the whole matrix $\mathbf{R}$ remains orthogonal. We find that this technique helps $R$ achieve lower training loss and produce a more ideal $\mathbf{R}\mathbf{W}$. In other words, we have a better chance to find global minimum. But we also find this becomes less effective when number of features and dimension become larger. This also calls for more study on this atypical optimization problem, i.e., with orthogonality constraint and ignorance of distance change to all but the nearest neighbor when computing gradients.
@@
We train $\mathbf{R}$ for toy models whose orthogonal structure is clear. Concretely, we present 4 experiments here, corresponding to toy models visualized in Fig. 1(a), Fig. 8(b), Fig. 9(a) and Fig. 9(b). The results are shown in Fig. 10, 11, 12, 13. We can see the proposed method is quite effective. Before merging, each subspace gradually learns to encode only one feature group. For example, in Fig. 10c, each 2D subspace in $\boldsymbol{\hat{h}}$ is computed either from the first or the second feature group (e.g., the first 2 rows only have nonzero entries in first 20 columns, corresponding to the first group). In Fig. 11, at the beginning the third subspace encodes all 4 groups (a), then after 3k steps it only encodes the 1st and 3rd group (b), then later it encodes only the 1st group. We can clearly see how NDM training objective leads $\mathbf{R}$ to evolve towards the desirable direction. Loss values also confirm the alignment between neighbor distances and “purity” of subspace (encoding one group instead of a mixture of groups). We denote the loss when search number is 4 as $l_{N=4}$, and similarly $l_{N=1024}$. We take the mean of loss across subspaces instead of sum when reporting them. For example, Fig. 10b’s loss values are: $l_{N=4}=0.212$ and $l_{N=1024}=0.011$. After more training when subspaces are “pure” (subfigure c), loss values are: $l_{N=4}=0.194$ and $l_{N=1024}=0.004$. Another example, Fig. 11 loss values are: $l_{N=4}=0.340$ and $l_{N=1024}=0.016$ for subfigure b, and $l_{N=4}=0.322$ and $l_{N=1024}=0.004$ for subfigure c. Importantly, we also see that $\mathbf{R}$ sometimes does not find global minimum. In Fig. 12 and 13, we see some sub-optimal subspaces, such as the 3rd subspace from the last in Fig. 12c, which is the reason for later imperfection of final $\mathbf{R}$. Moreover, we also see that the proposed MI-based merging works well. It correctly merges the subspaces encoding the same group. For example, in Fig. 11, normalized MI between 1st and 5th subspace is 0.55, and between 2nd and 6th is 0.49, normalized MI between all other pairs is below 0.03. In Fig. 10 and 11, we get the correct number of subspace and their dimension. In Fig. 12 and 13, the final configuration is imperfect but close to the ground truth. As mentioned before, we generally find the imperfection results from $\mathbf{R}$ being unable to find global minimum or disentangle feature groups for a few subspaces, the MI estimate usually works well.
@@
In summary, we find the training objective of NDM indeed encourages subspaces to encode only one feature group. Together with MI based merging, the final subspace partition is close to the groundtruth, with each subspace encoding one feature group. While training objective aligns well with our goal, the orthogonal matrix sometimes gets stuck at local minimum of this objective, resulting in sub-optimal partition. Nevertheless, it might not be necessary to achieve perfect partition to make itself useful in practice. In subfigure b of Fig. 10-13, we observe that these sub-optimal points already have a good deal of disentanglement, which could potentially be very useful when using these subspaces as analyzing unit instead of random subspaces.
@@ D.2 GPT-2 Small Test Suite
The Indirect Object Identification (IOI) task consists of sentences like: “When Mary and John went to the store, John gave a drink to _”, which should be completed with “Mary”. (Wang et al., 2022) found a circuit in GPT-2 Small responsible for solving this task (Fig. 14). Besides computational graph, they also reveal the information moved between the nodes of the graph.
Figure 14: IOI circuit, Fig. 2 from (Wang et al., 2022). Attention heads in this figure is denoted as layer_index.head_index. E.g., The second previous token head, 4.11, refers to layer 4 and attention head 11 in that layer. Note that indices start from 0.
To understand our intervention, we briefly describe the information flow inside the circuit: At token S1+1 (see Fig. 14 for the notation), the model attends the previous token and moves the name from it, which will be used as key in later induction heads. Right after layer 4, the model finishes moving previous token information, and this information is not yet used by downstream components. At token S2, the induction heads attend to S1+1, moving the relative position of S1 (i.e., it occurs before or after IO, or equivalently, S1+1 is “and” or “went”). Meanwhile, Duplicate token heads attend to S1 from S2, moving S name and relative position of S1. Therefore, at S2’s residual stream, right after layer 6 the model finishes moving position information of S1, and it is not yet moved by any downstream component. Then at END token, S-Inhibition heads attend to S2 and move the position information of S1 and as well as the S name. Right after layer 8, the model finishes moving position information of S1 and information of S name, and they are not yet used by any name mover heads. Lastly, name mover heads avoid attending S1 by using those information and attend to IO to move IO name. Note that the claims about the moved information is supported by strong causal evidence from (Wang et al., 2022), and is confirmed again with a different method in (Huang et al., 2024b).
@@
The Greater-than task (Hanna et al., 2023) consists of sentences in the form “The <noun> lasted from the year XXYY to the year XX”, such as “The war lasted from the year 1732 to the year 17”, the model should assign higher probability to years greater than 32, thus performing the simple greater-than math operation. (Hanna et al., 2023) revealed a circuit in GPT-2 Small responsible for solving this task. Briefly speaking, some attention heads between layer 5 to layer 9 (both inclusive) move YY to the last token’s residual stream, and MLP layer 8 to 11 (both inclusive) specify which years are greater than YY. Therefore, right after layer 9, the model finishes moving YY to the last token, so we intervene at $\boldsymbol{h}_{\text{END}}^{9,\text{post}}$ to change the YY, where END denotes the last token.
@@ D.3 NDM Training Details on GPT-2 Test Suite
For reproducibility, we report the hyperparameters used as follows: We use euclidean distance as the optimization objective. We use 32 as the unit size. We try different search numbers $N=\{1\times 2^{14},5\times 2^{14},25\times 2^{14}\}$. We do not do random re-initialization during the pre-merging stage like we did for the toy setting; preliminary results show this is not helpful for language models. We use batch size 128, learning rate 0.0003, and Adam optimizer. We try different merging intervals, 4k and 8k, and set the merging start delay as 2.5 times the merging interval. Longer merging intervals mean more training steps in general. We set 100k as the maximum training steps, which is usually not reached. We also try different merging threshold $\{0.02,0.03,0.04,0.05\}$. We use MiniPile (Kaddour, 2023) as the data to run GPT-2 Small when collecting activations, instead of the OpenWebText dataset which is similar to GPT-2 pretraining data. Therefore, we do hyperparameter search over all the combinations of different search numbers, merging interval, and merging threshold, while other hyperparameters are always the same. For each hyperparameter configuration, we train 4 orthogonal matrices for post-MLP residual stream at layer 4, 6, 8, 9 respectively, as they are needed in the test suite.
@@ D.4 Representing Patching Effect Distribution with Gini Coefficient
$\text{Var}_{s}$ ($\times 10^{3}$) 2.1 2.7 1.3 0.6 0.8 1.4 1.7 1.0 1.0 0.4 4.0 4.8 0.3
$\Delta P_{s}$ ($\times 10^{-2}$) 5.4 0.7 1.8 1.9 29.7 0.6 0.6 0.3 0.6 0.8 1.2 0.7 0.7
$\frac{\Delta P_{s}}{d_{s}}$ ($\times 10^{-4}$) 4.2 0.8 1.9 3.0 46.4 1.0 0.9 0.9 2.0 2.4 3.9 2.0 2.3
$\frac{\Delta P_{s}}{\text{Var}_{s}}$ ($\times 10^{-5}$) 2.6 0.3 1.4 2.9 39.6 0.5 0.3 0.3 0.6 2.2 0.3 0.1 2.2
@@
$\text{Var}_{s}$ ($\times 10^{3}$) 0.3 0.3 0.2 0.2 0.3 0.3 0.4 0.3 0.3 0.4 10.7
$\Delta LD_{s}$ ($\times 10^{-1}$) 14.8 11.6 7.6 4.2 3.1 2.5 2.9 1.4 1.6 0.9 2.2
$\frac{\Delta LD_{s}}{d_{s}}$ ($\times 10^{-3}$) 7.7 9.1 11.9 6.6 4.8 3.9 4.6 4.4 5.0 2.9 6.8
$\frac{\Delta LD_{s}}{\text{Var}_{s}}$ ($\times 10^{-4}$) 49.5 37.6 38.5 18.1 11.0 7.5 7.1 5.5 5.2 2.2 0.2
@@
$\text{Var}_{s}$ ($\times 10^{3}$) 0.7 1.0 0.5 7.9 0.4 0.2 0.2 0.1 0.1 0.1 0.1
$\Delta LD_{s}$ ($\times 10^{-1}$) 3.9 3.7 2.3 2.5 1.6 1.7 1.7 0.7 0.8 0.8 0.7
$\frac{\Delta LD_{s}}{d_{s}}$ ($\times 10^{-3}$) 3.0 2.9 2.4 2.6 2.4 2.6 2.7 2.2 2.4 2.4 2.2
$\frac{\Delta LD_{s}}{\text{Var}_{s}}$ ($\times 10^{-4}$) 5.5 3.6 4.7 0.3 4.0 8.1 9.1 5.2 8.0 7.7 7.0
@@ D.5 Ablation Study on GPT-2 Test Suite
search number ($\times 2^{14}$) merging threshold merging interval number of subspaces avg coefficient
1 0.02 4000 6,8,6,5 0.55
1 0.02 8000 9,9,9,9 0.64
1 0.03 4000 9,11,11,10 0.66
1 0.03 8000 9,11,11,10 0.61
1 0.04 4000 12,13,12,14 0.57
1 0.04 8000 12,12,13,14 0.63
1 0.05 4000 14,16,17,19 0.52
1 0.05 8000 13, 17,18,18 0.58
5 0.02 4000 5,8,7,6 0.58
5 0.02 8000 9,9,9,9 0.63
5 0.03 4000 8,8,9,10 0.64
5 0.03 8000 9,9,9,9 0.64
5 0.04 4000 11, 12, 14, 14 0.65
5 0.04 8000 10, 12, 13, 13 0.65
5 0.05 4000 12, 15, 16, 17 0.61
5 0.05 8000 12, 15, 16, 16 0.63
25 0.02 4000 5, 6, 5, 6 0.57
25 0.02 8000 9,9,9,9 0.66
25 0.03 4000 7, 9, 8, 10 0.63
25 0.03 8000 9, 9, 9, 10 0.65
25 0.04 4000 10, 10, 12, 13 0.68
25 0.04 8000 11, 11, 12, 13 0.71
25 0.05 4000 10, 13, 15, 16 0.63
25 0.05 8000 12, 15, 14, 15 0.71
Table 5: Combinations of different search number, merging threshold, and merging interval, and their resulting number of subspaces in the 4 layers used in the test (from low to high) and resulting average Gini coefficient on GPT-2 test suite. The one shown in Table 1 is the third row from the last, which is slightly better than the last row if we show more precision. Other hyperparameters are fixed, and described in D.3.
@@
search number ($\times 2^{14}$) merging threshold avg coefficient
@@
search number ($\times 2^{14}$) merge threshold avg coefficient
@@ D.6 Knowledge Conflict Experiment
Hyperparameters for NDM are the same as those for GPT-2 (App. D.3), except for the following ones: We use 64 as the unit size for Qwen2.5-1.5B and 72 for Gemma-2-2B. We use $N=25\times 2^{14}$ as the search number. We set 150k as maximum training steps. We try different merging threshold $\{0.015,0.02,0.03,0.04\}$ for Qwen and add one additional value 0.01 for Gemma. We try all these combinations and choose the best one. For Qwen2.5-1.5B, it is threshold = 0.02, merging interval = 8000; FOr Gemma-2-2B, it is threshold = 0.01, merging interval = 4000; We apply NDM to $\boldsymbol{h}^{11,mid}$ of Qwen2.5-1.5B (totally 28 layers) and $\boldsymbol{h}^{9,mid}$ of Gemma-2-2B (totally 26 layers). We choose the layer according to preliminary experiments where we patch the whole activation at the last token position of target entity.
@@
After training the estimated MI decreases significantly, see Fig. 17 and 18. Interestingly, as we can see, the original basis of the representation space of Qwen2.5-1.5B is quite special, simply partition on standard basis can produce many independent subspaces.
@@
We show baseline results in Fig. 19 and 20.